OpenStudy (anonymous):

verify cot^2t/csct=csct-sin^2

OpenStudy (anonymous):

I'm going to ignore all the t's, just see if you can follow along. (cos^2/sin^2)*sin=cos^2/sin cos^2=1-sin^2 therefore (1-sin^2)/sin=(1/sin)-(sin^2/sin)=csc-sin I'm not sure how you're getting the sin^2, cause all my math checks out, so that is pretty odd